Show that the sequence (an), where an= n/n^2+4 is monotonic. Is (an) a Cauchy sequence? Justify your answer


1
Expert's answer
2021-03-31T11:54:10-0400

"\displaystyle a_n = \frac{n}{n^2+4}"

The sequence is monotonic if "a_{n+1} < a_n \Rightarrow (a_{n+1}-a_n) <0".

"\displaystyle a_{n+1}=\frac{n+1}{(n+1)^2+4}=\frac{n+1}{n^2 +2n+5}"

"\displaystyle \frac{n+1}{n^2+2n+5} - \frac{n}{n^2+4} = \frac{(n+1)(n^2+4)-(n^2+2n+5)n}{(n^2+2n+5)(n^2+4)} = \frac{n^3+4n+n^2+4 -n^3-2n^2-5n}{(n^2+2n+5)(n^2+4)} = \frac{-n^2-n+4}{(n^2+2n+5)(n^2+4)}"

"-(n^2+n-4) <0" is true for any "n\geq 2".

Nevertheless, this sequence is not monotonic because "a_2>a_1".

This sequence is Cauchy sequence because there exists such a "n=n_0" that for all "n>n_0, p>n_0" and any "\epsilon>0:" "| a_{n+p} -a_n| < \epsilon" . Starting from n=2 this indeed works for any "\epsilon>0".